LUCIFER/LBC Technical Observing, prep for Partner Science:
UT 2010 Nov 01

Observer: DThompson, OKuhn in Tucson
Telescope Operator: GBechetti, SAllanson
Instrument Support: DOfficer
Telescope Support: MMidkiff
SW support: SW Build #AO5

Summary

Details

Afternoon: Red hub moved +1mm (away from M1) based on prior night's results. Telescope configured back to LBCs but there is a problem with a pump on the dynamic balance. As of 01:45 DAshby still on the phone trying to get things fixed.

01:45 Winds below 10m/s. T=5.7, RH=10 mostly clear.

Data: luci.20101101.NNNN.fits

01:59 LUCIFER darks (use LUCIFER$/readnoise script)
  • O2DCR 2.0s (0006-0007) 11.55 e-
  • MER10 10.0s (0008-0009) 6.47 e- (some light leakage)
  • Lights off in dome, blind+blind filters but no mask in FPU.

02:16 Dome open

LBC:

02:19 WT10_336 to collimate at high EL. Red channel filter hardware filter error. Trying dofpia anyway....

02:22 dofpia, /first - nope, still a problem. Stopped dofpia with "S" backed out the initial Z11s sent.

02:25 shutdown and restart LBC from power control. Everything came back up. Olga says this appears to be a communication problem and not a stuck filter wheel like before.

02:30 dofpia, /first (failed after first focus offset. May need to add coma by hand)

02:35 dofpia (failed again) Coma offsets: RED offset Z7 -2000nm, BLUE offset Z7 -2000nm

02:43 dofpia (failed again) Red -1000 Z8, BLUE +100 Z22 (pupils have a peak in the middle)

03:00 We hit the -Z limit on red again, almost exactly 1 mm higher than yesterday. Pupils are almost collimated. Doug is willing to bump red another millimeter so we are closing the dome to get this done.

Still diverging, but now it looks like Z11 rather than Z22. Clear c00 forces. Exit IDL and cd to /home/LBTO and try again

05:04 dofpia, /first, /redonly Looked OK for a while but still diverging. Olga and Andrew will try some things.

Test lbcrangebal.cl

Take a set of Collimation data

11:48 Conclusion is that one of the two filter wheels is still out of alignment and when it should have a "clear" in place it has a red filter.

11:48 Switching to LUCIFER+IRTC
12:20 M3 did not complete its motion.

LUCIFER:

Mask configuration GUI updated with new masks. Checking:
Mask Luci# Mask ID   Mask Luci# Mask ID
A0068 0010 903405   MPEQ2343a 0017 931447
CL0128M6 0011 932651   MPEQ2343b 0018 931721
CL0218M4 0012 915954   MPETELL   9
CL0218M5 0014 925143   RDSEC   9
Cl0016 0015 902928   Wasp12b   9
LSWC03a 0016 980100   mclinden   9

LBC + Telescope problems meant we never went on-sky with LUCIFER.

-- DavidThompson - 01 Nov 2010
Topic revision: r4 - 01 Nov 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback